# Semantic Web: Past, Present, and Future

### v1

@@ Abstract
In this paper, we recap the classical concepts and foundations of the Semantic Web as well as modern and recent concepts and applications, building upon these foundations. The classical topics we cover include knowledge representation, creating and validating knowledge on the Web, reasoning and linking, and distributed querying. We enhance this classical view of the so-called “Semantic Web Layer Cake” with an update of recent concepts that include provenance, security and trust, as well as a discussion of practical impacts from industry-led contributions. We conclude with an outlook on the future directions of the Semantic Web.
@@
This article is a living document. The first version was written by Jannik, Scherp, and Staab in 2011 [95]. It was translated to German and updated by Gröner, Scherp, and Staab in 2013 [72] and updated again by Gröner and Scherp in 2020 [141]. This release in 2024 reflects on the latest developments in knowledge graphs and large language models. It has been translated back to English and extended with contributions by Petr Škoda, Katja Hose, and Maria-Esther Vidal.
@@ 1 Introduction
The vision of the Semantic Web as coined by Tim Berners-Lee, James Hendler, and Orla Lassila [22] in 2001 is to develop intelligent agents that can automatically gather semantic information from distributed sources accessible over the Web, integrate that knowledge, use automated reasoning [69], and solve complex tasks as such as schedule appointments in negotiation of the preferences of the involved parties. We have come a long way since then. In this paper, we reflect on the past, i. e., the ideas and components developed in the early days of the Semantic Web. Since the beginning, the Semantic Web has tremendously developed and undergone multiple waves of innovation. The Linked Data movement has especially seen uptake by industries, governments, and non-profit organizations, alike. We discuss those present components and concepts that have been added over the years and shown to be very useful. Although many concepts of the initial idea of the Semantic Web have been implemented and put into practice, still further research is needed to reach the full vision. Thus, this paper concludes with an outlook to future directions and steps that may be taken.
@@
For the novice reader of the Semantic Web, we provide a brief historical overview of the developments and innovation waves of the Semantic Web: At the beginning of the Semantic Web, we were mainly talking about publishing Linked Data on the Web [78], i. e., semantic data typically structured using the Resource Description Framework (RDF) [^1] http://www.w3.org/TR/rdf-primer/ that is accessible on the Web using URIs/IRIs to identify entities, classes, predicates, etc. By referencing entities from other websites and Web-accessible sources, i. e., dereferencable via HTTP, the data becomes naturally linked. By using standardized vocabularies and ontologies the information then becomes more aligned and easier to use across sources. These principles have allowed non-profit organizations, companies, governments, and individuals to publish and share large amounts of interlinked data, which has led to the success of the Linked Open Data cloud [^2] https://lod-cloud.net/ since 2007. Since many of the large interconnected semantic sources are accessible via interfaces understanding structured query languages (SPARQL endpoints), federated query processing methods were developed that allow exploiting the strengths of structured query languages to precisely formulate an information need and optimize the query for efficient execution in a distributed setting.
@@
When Google launched its Knowledge Graph in 2012333https://blog.google/products/search/introducing-knowledge-graph-things-not/, semantic technologies experienced another wave of new applications in the context of searching information. Whereas search engines before mainly relied on keyword search and string-based matches of the keywords in the websites’ text, the knowledge graph enabled including semantics to capture the user’s information need as well as the meaning of potentially relevant documents. To achieve this purpose, Google’s knowledge graph integrates large amounts of machine-processable data available on the Web and uses this information not only to improve search results but also to display infoboxes for entities identified in the user’s keywords. It is only since 2012 that we have widely used the term “knowledge graph” to refer to semantic data, where entities are connected via relationships and form large graphs of interconnected information, typically with RDF as a common standard language. In recent years though (labeled) property graphs (LPG) have been used to manage knowledge graphs. We refer to the literature for a detailed comparison of RDF graphs and LPGs [85] and also like to point out that they can be converted into each other [26]. In this article, we consider knowledge graphs from the perspective of the Semantic Web, i. e., we consider RDF graphs.
@@
A few years later, semantic technologies found another novel application in enterprise settings as enterprise knowledge graphs [60, 119] and data integration [36, 170]. Since all kinds of information can be structured as a graph, knowledge graphs can be used as a common structure to integrate heterogeneous information that is otherwise locked up in silos in different branches of a company. Integrating the data into a knowledge graph then allows for an integrated view, efficiently retrieving relevant information from this view once needed, and integrating external information that is available in the form of knowledge graphs, for instance on the Semantic Web. After all, online analytical processing-style queries can also be formulated with SPARQL [^4] http://www.w3.org/TR/sparql11-query/ and evaluated on (distributed) knowledge graphs.
@@
In the past few years, learning on graph data became one of the fastest growing and most active areas in machine learning [76]. Graph representation learning has created a new wave of graph embedding models and graph neural networks on knowledge graphs for tasks such as entity classification and link prediction. Natural Language Processing (NLP) has been another important field of the Semantic Web since the early years to extract knowledge from textual data and make it machine readable. Another field where NLP meets the Semantic Web is user interfaces for search on structured data to enable intuitive, natural language querying for graph data [74] similar to web search engines. At the end of 2022, ChatGPT [^5] https://chat.openai.com/ emerged as the first publicly available end-consumer tool based on a Large Language Model (LLM). Since then, the GPT-based family of LLMs has stirred up research and business alike and demonstrated impressive performance on many NLP tasks, including generating structured queries from user prompts and extracting structured knowledge from text [156].
@@
In this paper, we provide a comprehensive overview of the Semantic Web with its semantic technologies and underlying principles that have been inspiring and driving the multiple waves of innovations in the past two decades. Section 2 provides a motivating example for the classic Semantic Web. We refer back to this example throughout the paper. Section 3 presents the principles and the general architecture of the Semantic Web along with the basic semantic technologies it is founded upon. Besides classical components, we are also describing recent developments, and pointing out components that are still being researched and developed. Section 4 shows how to represent distributed knowledge on the Semantic Web. The creation and maintenance of graph data is described in Section 5. Section 6 discusses the principle of reasoning and logical inference. Section 7 then shows how to query over the (distributed) graph data on the Semantic Web. We discuss the trustworthiness and provenance of data on the Semantic Web in Section 8. We provide extensive examples of applications based on and using the Semantic Web and its technologies in Section 9. Finally, we reflect on the impact the Semantic Web has on practitioners in Section 10. Finally, we conclude with a brief outlook on future developments for the Semantic Web.
@@ 2 Motivating Example
We consider the example of the BBC program ontology with links to various other ontologies such as for music, events, and social networks as shown in Figure 1. We start with the BBC playlists of its radio stations. The playlists are published online in Semantic Web formats. We can leverage the playlist to get unique identifiers of played artists and bands. For example, the music group “ABBA” has a unique identifier in the form of a URI (https://www.bbc.co.uk/programmes/b03lyzpr). This URI can be used to link the music group to information from the MusicBrainz [^9] http://musicbrainz.org/ music portal. MusicBrainz knows the members of the band, such as Benny Andersson, as well as the genre and songs. In addition, MusicBrainz is linked to Wikipedia [^10] https://www.wikipedia.org/ (not shown in the figure), e. g., to provide information about artists, such as biographies on DBpedia [12]. Information about British radio stations can be found in the form of lists on Web pages such as Radio UK [^11] https://www.radio-uk.co.uk, which can also be converted into a representation in the Semantic Web.
@@
Colloquially, an ontology is a description of concepts and their relationships. Ontologies are used to formally represent knowledge on the Semantic Web. [^12] An ontology definition is provided in Section 4.2. For example, Dublin Core [^13] http://dublincore.org/documents/dc-rdf/ provides a metadata schema for describing common properties of objects, such as the creator of the information, type, date, title, usage rights, and so on. Figure 1 presents ontologies used to describe data in our example. For example, the Playcount ontology [^14] http://dbtune.org/bbc/playcount/ of the BBC is used to model which artist was played and how many times in the programs. Ontologies can be interconnected in the Semantic Web. For example, the MusicBrainz ontology is connected to the BBC ontology using the Playcount ontology. Different ontologies with varying degrees of formality and different relationships to each other are used by the BBC to describe their data (see also [133]).
@@ 3 Architecture of the Semantic Web
The example in Section 2 describes what the Semantic Web is as an infrastructure, but not how this is achieved. In fact, the capabilities of the Semantic Web in a small scale have already been implemented by some knowledge-based systems originating from artificial intelligence research, e. g., Heinsohn et al [82]. However, for the implementation of the vision on a large scale, i. e, the Web, these knowledge-based systems lacked flexibility, robustness, and scalability. In part, this was due to the complexity of the algorithms used. For example, knowledge bases in description logic in the 1990s, which serve as the basis of Web ontologies, were limited regarding their size such that they could handle at the most some hundred concepts [82].
@@
Entities (also called resources) are identified on the Internet by so-called Uniform Resource Identifiers (URIs) [20]. When a URI holds a dereferenceable location of the resource, in other words, it can be employed to get access to the resource via HTTP, it is called a Uniform Resource Locator (URL) [23, 21]. Furthermore, Internationalized Resource Identifiers (IRIs) [49] supplement URIs with international character sets from Unicode/ISO10646. URIs are globally and universally used but are usually not under our control. A recent W3C recommendation, the Decentralized Identifiers (DIDs), introduces an alternative approach to the above identifiers [148]. A DID is by default decentralized and allows for self-sovereign management of the identity, i. e., the control of a DID and its associated data is with the users.
@@
In addition to the referencing of resources and a uniform syntax for the exchange of data, a data model is required that allows resources to be described both individually and in their entirety and how they are linked [78, 81]. An integrated representation of data from multiple sources is provided by a data model based on directed graphs [125]. The corresponding W3C standard language is RDF (Resource Description Framework) [^26] https://www.w3.org/RDF/.
An RDF graph consists of a set of RDF triples, where a triple consists of a subject, predicate (property), and object. An RDF graph can be given an identifier, such a graph is called a named graph [32]. RDF graphs can be serialized in several ways (see Syntax for Data Exchange above). It is important to note that some formats (e. g., Turtle) do not support named graphs. Finally, RDF-star (aka RDF*) [^27] https://w3c.github.io/rdf-star/, [^28] https://blog.liu.se/olafhartig/2019/01/10/position-statement-rdf-star-and-sparql-star/ was introduced to allow nesting of triples and thus enables an efficient way to make statements about statements while avoiding reification and the increased number of triples and complexity that come along with it.
@@
Validating constraints, representing syntactic and semantic restrictions in RDF graphs, is essential for ensuring data quality. In addition to rule-based languages, shapes allow for the specification of conditions to meet data quality criteria and integrity constraints. A shape encompasses a conjunction of constraints representing conditions that nodes in an RDF graph must satisfy [86]. A shapes graph is a labeled directed graph where nodes correspond to shapes, and edges denote interrelated constraints. The Shapes Constraint Language (SHACL) [103] and Shape Expressions (ShEx) [128]) are two W3C-recommendations to express shapes graphs over RDF [130].
@@
RDFS is not expressive enough to merge data from different sources and define consistency criteria about it, such as the disjointness of classes or the equivalence of resources. The Web Ontology Language (OWL) [^34] https://www.w3.org/OWL/ [1] is an ontology language with formally defined meaning based on description logic. This allows for reasoning services to be provided by knowledge-based systems for OWL ontologies. OWL can be exchanged using RDF data formats. Compared to RDFS, OWL provides more expressive language constructs. For example, OWL allows the specification of equivalences between classes and cardinality constraints on properties [1].
@@
Finally, there is RDF-star (aka RDF*) [^37] https://w3c.github.io/rdf-star/, [^38] https://blog.liu.se/olafhartig/2019/01/10/position-statement-rdf-star-and-sparql-star/ – along with the corresponding SPARQL-star/ SPARQL* extension – was proposed and since then was implemented by several triple stores [3] that often provide publicly accessible SPARQL endpoints. The key idea with RDF/SPARQL-star is to allow the nesting of triples to enable an efficient way to allow statements about statements while avoiding reification and the increased number of triples and complexity that come along with it.
@@
Other aspects of the Semantic Web are encryption and authentication to ensure that data transmissions cannot be intercepted, read, or modified. Crypto modules, such as SSL (Secure Socket Layer), verify digital certificates and enable data protection and authentication. In addition, there are digital signatures for graphs that integrate seamlessly into the architecture of the Semantic Web and are themselves modeled as graphs again [100]. This allows graph signatures to be applied iteratively and enables to building trust networks. The Verifiable Credentials Data Model, a recent W3C recommendation, introduces a standard to model trustworthy credentials for graphs on the web [^39] https://www.w3.org/TR/vc-data-model/. Data on the Semantic Web can be augmented with additional information about its trustworthiness and provenance.

### v2

@@ Abstract
In this paper, we recap the classical concepts and foundations of the Semantic Web as well as modern and recent concepts and applications, building upon these foundations. The classical topics we cover include knowledge representation, creating and validating knowledge on the Web, reasoning and linking, and distributed querying. We enhance this classical view of the so-called “Semantic Web Layer Cake” with an update of recent concepts. These include provenance, security and trust, as well as a discussion of practical impacts from industry-led contributions. We also provide an overiew of shallow and deep machine learning methods for knowledge graphs and discuss the relation of language models and knowledge graphs. We conclude with an outlook on the future directions of the Semantic Web.
@@
Changelog
  • Update in November 2025: Added an overview of machine learning methods for knowledge graphs in Section 9, particularly shallow graph embeddings and neural networks for graphs. We discuss language models and their relation to and use for knowledge graphs in Section 10.
@@
This article is a living document. The first version was written by Jannik, Scherp, and Staab in 2011 [DBLP:journals/insk/JanikSS11]. It was translated to German and updated by Gröner, Scherp, and Staab in 2013 [DBLP:books/ol/13/GronerSS13] and updated again by Gröner and Scherp in 2020 [DBLP:books/degruyter/20/ScherpG20]. This release in 2024 reflects on the latest developments in knowledge graphs and large language models. It has been translated back to English and extended with contributions by Petr Škoda, Katja Hose, and Maria-Esther Vidal.
@@ 1 Introduction
The vision of the Semantic Web as coined by Tim Berners-Lee, James Hendler, and Orla Lassila [TimBernersLee2001] in 2001 is to develop intelligent agents that can automatically gather semantic information from distributed sources accessible over the Web, integrate that knowledge, use automated reasoning [DBLP:journals/ki/GlimmS16a], and solve complex tasks as such as schedule appointments in negotiation of the preferences of the involved parties. We have come a long way since then. In this paper, we reflect on the past, i. e., the ideas and components developed in the early days of the Semantic Web. Since the beginning, the Semantic Web has tremendously developed and undergone multiple waves of innovation. The Linked Data movement has especially seen uptake by industries, governments, and non-profit organizations, alike. We discuss those present components and concepts that have been added over the years and shown to be very useful. Although many concepts of the initial idea of the Semantic Web have been implemented and put into practice, still further research is needed to reach the full vision. Thus, this paper concludes with an outlook to future directions and steps that may be taken.
@@
For the novice reader of the Semantic Web, we provide a brief historical overview of the developments and innovation waves of the Semantic Web: At the beginning of the Semantic Web, we were mainly talking about publishing Linked Data on the Web [DBLP:books/crc/linked2014], i. e., semantic data typically structured using the Resource Description Framework (RDF) [^1] http://www.w3.org/TR/rdf-primer/ that is accessible on the Web using URIs/IRIs to identify entities, classes, predicates, etc. By referencing entities from other websites and Web-accessible sources, i. e., dereferencable via HTTP, the data becomes naturally linked. By using standardized vocabularies and ontologies the information then becomes more aligned and easier to use across sources. These principles have allowed non-profit organizations, companies, governments, and individuals to publish and share large amounts of interlinked data, which has led to the success of the Linked Open Data cloud [^2] https://lod-cloud.net/ since 2007. Since many of the large interconnected semantic sources are accessible via interfaces understanding structured query languages (SPARQL endpoints), federated query processing methods were developed that allow exploiting the strengths of structured query languages to precisely formulate an information need and optimize the query for efficient execution in a distributed setting.
@@
When Google launched its Knowledge Graph in 2012333https://blog.google/products/search/introducing-knowledge-graph-things-not/, semantic technologies experienced another wave of new applications in the context of searching information. Whereas search engines before mainly relied on keyword search and string-based matches of the keywords in the websites’ text, the knowledge graph enabled including semantics to capture the user’s information need as well as the meaning of potentially relevant documents. To achieve this purpose, Google’s knowledge graph integrates large amounts of machine-processable data available on the Web and uses this information not only to improve search results but also to display infoboxes for entities identified in the user’s keywords. It is only since 2012 that we have widely used the term “knowledge graph” to refer to semantic data, where entities are connected via relationships and form large graphs of interconnected information, typically with RDF as a common standard language. In recent years though (labeled) property graphs (LPG) have been used to manage knowledge graphs. We refer to the literature for a detailed comparison of RDF graphs and LPGs [DBLP:journals/corr/abs-2003-02320] and also like to point out that they can be converted into each other [DBLP:phd/dnb/Blume22]. In this article, we consider knowledge graphs from the perspective of the Semantic Web, i. e., we consider RDF graphs.
@@
A few years later, semantic technologies found another novel application in enterprise settings as enterprise knowledge graphs [DBLP:conf/iceis/GalkinAVS17, Noy19] and data integration [Cheatham2017, DBLP:books/sp/10/Wood10]. Since all kinds of information can be structured as a graph, knowledge graphs can be used as a common structure to integrate heterogeneous information that is otherwise locked up in silos in different branches of a company. Integrating the data into a knowledge graph then allows for an integrated view, efficiently retrieving relevant information from this view once needed, and integrating external information that is available in the form of knowledge graphs, for instance on the Semantic Web. After all, online analytical processing-style queries can also be formulated with SPARQL [^4] http://www.w3.org/TR/sparql11-query/ and evaluated on (distributed) knowledge graphs.
@@
In the past few years, learning on graph data became one of the fastest growing and most active areas in machine learning [DBLP:series/synthesis/2020Hamilton]. Graph representation learning has created a new wave of graph embedding models and graph neural networks on knowledge graphs for tasks such as entity classification and link prediction. Natural Language Processing (NLP) has been another important field of the Semantic Web since the early years to extract knowledge from textual data and make it machine readable. Another field where NLP meets the Semantic Web is user interfaces for search on structured data to enable intuitive, natural language querying for graph data [DBLP:journals/eswa/HabernalK13] similar to web search engines. At the end of 2022, ChatGPT [^5] https://chat.openai.com/ emerged as the first publicly available end-consumer tool based on a Large Language Model (LLM). Since then, the GPT-based family of LLMs has stirred up research and business alike and demonstrated impressive performance on many NLP tasks, including generating structured queries from user prompts and extracting structured knowledge from text [DBLP:journals/corr/abs-2305-04676].
@@
In this paper, we provide a comprehensive overview of the Semantic Web with its semantic technologies and underlying principles that have been inspiring and driving the multiple waves of innovations in the past two decades. Section 2 provides a motivating example for the classic Semantic Web. We refer back to this example throughout the paper. Section 3 presents the principles and the general architecture of the Semantic Web along with the basic semantic technologies it is founded upon. Besides classical components, we are also describing recent developments, and pointing out components that are still being researched and developed. Section 4 shows how to represent distributed knowledge on the Semantic Web. The creation and maintenance of graph data is described in Section 5. Section 6 discusses the principle of reasoning and logical inference. Section 7 then shows how to query over the (distributed) graph data on the Semantic Web. We discuss the trustworthiness and provenance of data on the Semantic Web in Section 8. Machine learning methods for knowledge graphs are discussed in Section 9, particularly shallow graph embeddings and neural networks for graphs. We discuss language models and their relation and use for knowledge graphs in Section 10. We provide extensive examples of applications based on and using the Semantic Web and its technologies in Section 11. Finally, we reflect on the impact the Semantic Web has on practitioners in Section 12. Finally, we conclude with a brief outlook on future developments for the Semantic Web.
@@ 2 Motivating Example
We consider the example of the BBC program ontology with links to various other ontologies such as for music, events, and social networks as shown in Figure 1. We start with the BBC playlists of its radio stations. The playlists are published online in Semantic Web formats. We can leverage the playlist to get unique identifiers of played artists and bands. For example, the music group “ABBA” has a unique identifier in the form of a URI (https://www.bbc.co.uk/programmes/b03lyzpr). This URI can be used to link the music group to information from the MusicBrainz [^9] http://musicbrainz.org/ music portal. MusicBrainz knows the members of the band, such as Benny Andersson, as well as the genre and songs. In addition, MusicBrainz is linked to Wikipedia [^10] https://www.wikipedia.org/ (not shown in the figure), e. g., to provide information about artists, such as biographies on DBpedia [key:dbpedia]. Information about British radio stations can be found in the form of lists on Web pages such as Radio UK [^11] https://www.radio-uk.co.uk, which can also be converted into a representation in the Semantic Web.
@@
Colloquially, an ontology is a description of concepts and their relationships. Ontologies are used to formally represent knowledge on the Semantic Web. [^12] An ontology definition is provided in Section 4.2. For example, Dublin Core [^13] https://www.dublincore.org/specifications/dublin-core/dc-rdf/ provides a metadata schema for describing common properties of objects, such as the creator of the information, type, date, title, usage rights, and so on. Figure 1 presents ontologies used to describe data in our example. For example, the Playcount ontology [^14] http://dbtune.org/bbc/playcount/ of the BBC is used to model which artist was played and how many times in the programs. Ontologies can be interconnected in the Semantic Web. For example, the MusicBrainz ontology is connected to the BBC ontology using the Playcount ontology. Different ontologies with varying degrees of formality and different relationships to each other are used by the BBC to describe their data (see also [Raimond2009IMR]).
@@ 3 Architecture of the Semantic Web
The example in Section 2 describes what the Semantic Web is as an infrastructure, but not how this is achieved. In fact, the capabilities of the Semantic Web in a small scale have already been implemented by some knowledge-based systems originating from artificial intelligence research, e. g., Heinsohn et al [DBLP:journals/ai/HeinsohnKNP94]. However, for the implementation of the vision on a large scale, i. e, the Web, these knowledge-based systems lacked flexibility, robustness, and scalability. In part, this was due to the complexity of the algorithms used. For example, knowledge bases in description logic in the 1990s, which serve as the basis of Web ontologies, were limited regarding their size such that they could handle at the most some hundred concepts [DBLP:journals/ai/HeinsohnKNP94].
@@
Entities (also called resources) are identified on the Internet by so-called Uniform Resource Identifiers (URIs) [rfc1630]. When a URI holds a dereferenceable location of the resource, in other words, it can be employed to get access to the resource via HTTP, it is called a Uniform Resource Locator (URL) [rfc1738, rfc3986]. Furthermore, Internationalized Resource Identifiers (IRIs) [rfc3987] supplement URIs with international character sets from Unicode/ISO10646. URIs are globally and universally used but are usually not under our control. A recent W3C recommendation, the Decentralized Identifiers (DIDs), introduces an alternative approach to the above identifiers [w3c-did]. A DID is by default decentralized and allows for self-sovereign management of the identity, i. e., the control of a DID and its associated data is with the users.
@@
In addition to the referencing of resources and a uniform syntax for the exchange of data, a data model is required that allows resources to be described both individually and in their entirety and how they are linked [DBLP:books/crc/linked2014, DBLP:reference/bdt/HartigHS19]. An integrated representation of data from multiple sources is provided by a data model based on directed graphs [key:oem]. The corresponding W3C standard language is RDF (Resource Description Framework) [^26] https://www.w3.org/RDF/.
An RDF graph consists of a set of RDF triples, where a triple consists of a subject, predicate (property), and object. An RDF graph can be given an identifier, such a graph is called a named graph [DBLP:conf/semweb/CarrollBHS04]. RDF graphs can be serialized in several ways (see Syntax for Data Exchange above). It is important to note that some formats (e. g., Turtle) do not support named graphs. Finally, RDF-star (aka RDF*) [^27] https://w3c.github.io/rdf-star/, [^28] https://blog.liu.se/olafhartig/2019/01/10/position-statement-rdf-star-and-sparql-star/ was introduced to allow nesting of triples and thus enables an efficient way to make statements about statements while avoiding reification and the increased number of triples and complexity that come along with it.
@@
Validating constraints, representing syntactic and semantic restrictions in RDF graphs, is essential for ensuring data quality. In addition to rule-based languages, shapes allow for the specification of conditions to meet data quality criteria and integrity constraints. A shape encompasses a conjunction of constraints representing conditions that nodes in an RDF graph must satisfy [DBLP:series/synthesis/2021Hogan]. A shapes graph is a labeled directed graph where nodes correspond to shapes, and edges denote interrelated constraints. The Shapes Constraint Language (SHACL) [SHACL] and Shape Expressions (ShEx) [SHEX]) are two W3C-recommendations to express shapes graphs over RDF [DBLP:conf/www/RabbaniLH22].
@@
RDFS is not expressive enough to merge data from different sources and define consistency criteria about it, such as the disjointness of classes or the equivalence of resources. The Web Ontology Language (OWL) [^34] https://www.w3.org/OWL/ [key:owl2] is an ontology language with formally defined meaning based on description logic. This allows for reasoning services to be provided by knowledge-based systems for OWL ontologies. OWL can be exchanged using RDF data formats. Compared to RDFS, OWL provides more expressive language constructs. For example, OWL allows the specification of equivalences between classes and cardinality constraints on properties [key:owl2].
@@
Finally, there is RDF-star (aka RDF*) [^37] https://w3c.github.io/rdf-star/, [^38] https://blog.liu.se/olafhartig/2019/01/10/position-statement-rdf-star-and-sparql-star/ – along with the corresponding SPARQL-star/ SPARQL* extension – was proposed and since then was implemented by several triple stores [AbuAebAgl23] that often provide publicly accessible SPARQL endpoints. The key idea with RDF/SPARQL-star is to allow the nesting of triples to enable an efficient way to allow statements about statements while avoiding reification and the increased number of triples and complexity that come along with it.
@@
Other aspects of the Semantic Web are encryption and authentication to ensure that data transmissions cannot be intercepted, read, or modified. Crypto modules, such as SSL (Secure Socket Layer), verify digital certificates and enable data protection and authentication. In addition, there are digital signatures for graphs that integrate seamlessly into the architecture of the Semantic Web and are themselves modeled as graphs again [DBLP:conf/esws/KastenSS14]. This allows graph signatures to be applied iteratively and enables to building trust networks. The Verifiable Credentials Data Model, a recent W3C recommendation, introduces a standard to model trustworthy credentials for graphs on the web [^39] https://www.w3.org/TR/vc-data-model/. Data on the Semantic Web can be augmented with additional information about its trustworthiness and provenance.